3g C and 5g oxygen react completely with each other to give both CO and CO2. Find mole fraction of CO2 in the resulting mixture. Please give the solution.

Grade:11

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Solution:
The reaction for the same would be given by:
3C+2O2\rightarrow 2CO+CO2
Now the number of moles of C are equal to:
\frac{3}{12}=\frac{1}{4}
The number of moles of Oxygen are:
\frac{5}{16}
Thus the limiting reagent is C.
And so the mole fraction of CO2 would be equal to:
\frac{\frac{1}{3}}{\frac{1}{3}+\frac{2}{3}}=\frac{1}{3}
Thanks & Regards
Sumit Majumdar,
askIITians Faculty
Ph.D,IIT Delhi
Sumit Majumdar IIT Delhi
9 years ago
Solution:
The reaction for the same would be given by:
3C+2O2\rightarrow 2CO+CO2
Now the number of moles of C are equal to:
\frac{3}{12}=\frac{1}{4}
The number of moles of Oxygen are:
\frac{5}{32}
So the limiting reagent is Oxygen.
And so the mole fraction of CO2 would be equal to:
\frac{\frac{1}{2}}{\frac{1}{2}+\frac{1}{1}}=\frac{1}{3}
Regards
Sumit